OpenStudy (anonymous):

5 volume of alcohol is mixed with 8 volume of water and how many volume of alcohol must be added so that the ratio of volume of alcohol to that of water is 5:1?

OpenStudy (anonymous):

\[\frac{ 5+x }{ 8 } = \frac{ 5 }{ 1 }\]

OpenStudy (anonymous):

solve it for it x

OpenStudy (anonymous):

35?

OpenStudy (anonymous):

yes that should be correct if solutions of water and alcohol keep the volume ratio

OpenStudy (anonymous):

thanks!